A boat can cruise at 25 mph in still water
What is its cruising speed upstream against a current of 7 mph?
What is its cruising speed downstream with a current of 7 mph?

Answer by srinivas.g(540)

speed of boat in still water = 25 mph
speed of current = 7 mph
speed of boat upstream against a current of 7 mph = 25-7 = 18 mph
speed of boat downstream with current of 7mph = 25+7 = 32 mph